OpenStudy (anonymous):

What is the solution of the system? y=9x-2 y=7x+3

OpenStudy (solomonzelman):

y=9x-2 y=7x+3 ------- 0y=2y-5 -2y=-5 y=5/2 Solve for x.

OpenStudy (anonymous):

Thank you.